OpenStudy (anonymous):

The probability of tossing heads with a weighted coin is 0.4. If this unfair coin is tossed 50 times, what is the probability that heads will appear between 15 and 25 times, inclusive? (Use the normal approximation, if valid.)

OpenStudy (anonymous):

p = 0.4 n = 50 mean= 20*0.4 = 8 > 5 (normal approximation valid.) sd= sqrt(npq) = sqrt(20*0.4*0.6) = 2.191 z = (15 - 8)/2.191 = 3.1948 z = (25 - 8)/2.191 = 7.7590 P(15 < x < 25) = P(3.1948 < z < 7.7590) = 0.0007

OpenStudy (anonymous):

is that true